# Improving Matrix Exponential for Generative AI Flows: A Taylor-Based Approach Beyond Paterson–Stockmeyer

### v1

@@ 5.1 Computational scaling and matrix dimensionality
This analysis evaluates how the performance of the proposed matrix exponential implementation, expm_flow_sastre, scales relative to the baseline, expm_flow, as matrix dimensionality increases. We examine matrix orders $n\in\{8,16,32,64,128,256,512,1024\}$ under two distinct scenarios: individual matrix transformations of size $n\times n$ and batched tensor operations of size $n\times 16\times 16$. The execution time required to compute 1000 matrix exponentials was recorded for each configuration, with the resulting trends illustrated in Figure 6.
@@
Figure 6: Comparison of the execution time required to calculate the matrix exponential using the original method (expm_flow) and the proposed method (expm_flow_sastre). The results are presented for different matrix sizes $n\in\{2,4,8,16,32,64,128,256,512,1024\}$. (Left) Performance analysis for single $n\times n$ matrices. (Right) Performance analysis for tensors of size $n\times 16\times 16$.
@@ 5.2 Training Performance and Scaling
To evaluate the practical utility of our approach, we integrated the proposed matrix exponential method into a generative framework. Glow [39] is a coupling layer approach to learning invertible transformations. We replaced the original implementation (expm_flow) with our optimized function (expm_flow_sastre) to measure its impact on computational efficiency during the training phase. All models were trained for 50 epochs across three benchmark datasets: CIFAR-10, ImageNet32, and ImageNet64. As detailed in Table 4, the proposed method demonstrates increased computational efficiency across all tested configurations. While the minimum observed speed-up was $3.91$ for ImageNet64, the most significant gain occurred with ImageNet32, where training was nearly 10 times faster ($9.74$) than the baseline. This dramatic reduction in training time, specifically the shift from 32,304 seconds to 3,317 seconds per epoch on ImageNet32, demonstrates that the computational bottleneck previously associated with matrix exponentials in flow-based models can be effectively mitigated.
@@
expm_flow_sastre time 443.4 3317 10875
@@
Table 4: Average training time per epoch (in seconds) for the matrix exponential version of Glow. Results compare the efficiency of our proposed expm_flow_sastre function against the original expm_flow implementation across three standard datasets.
@@ 5.3 Inference and Sampling Performance
While the previous sections evaluated performance during the training phase, efficiency during inference is equally critical for generative models. Table 5 presents a comparison of the sampling times between expm_flow and expm_flow_sastre. In this context, sampling time refers to the duration required for the model to generate an image from the latent space. We evaluated two scenarios: generating a single image ($n=1$) and generating a batch of 128 images ($n=128$). The results demonstrate that the speedup provided by our method becomes significantly more pronounced as the batch size increases. While the performance is nearly identical for a single sample, the expm_flow_sastre variant achieves a speedup factor of approximately $1.95$ for larger batches, effectively halving the inference latency.
@@
expm_flow_sastre time 0.3836 0.4247
@@
Table 5: Average inference times (in seconds) for sampling a matrix exponential variant of Glow. We compare the original expm_flow against expm_flow_sastre for single and batched samples, including the calculated speedup factor.
@@ 6 Conclusions
These results demonstrate that modern polynomial evaluation strategies allow for higher-order approximations at a lower computational cost, measured in terms of matrix multiplications ($M$), than traditional rational or Paterson–Stockmeyer-based methods. Consequently, the proposed implementation offers a portable, library-independent solution for high-throughput applications and large-scale simulation environments.

### v2

@@ 5.1 Computational scaling and matrix dimensionality
This analysis evaluates how the performance of the proposed matrix exponential implementation, expm_flow_opt15, scales relative to the baseline, expm_flow, as matrix dimensionality increases. We examine matrix orders $n\in\{8,16,32,64,128,256,512,1024\}$ under two distinct scenarios: individual matrix transformations of size $n\times n$ and batched tensor operations of size $n\times 16\times 16$. The execution time required to compute 1000 matrix exponentials was recorded for each configuration, with the resulting trends illustrated in Figure 6.
@@
Figure 6: Comparison of the execution time required to calculate the matrix exponential using the original method (expm_flow) and the proposed method (expm_flow_opt15). The results are presented for different matrix sizes $n\in\{2,4,8,16,32,64,128,256,512,1024\}$. (Left) Performance analysis for single $n\times n$ matrices. (Right) Performance analysis for tensors of size $n\times 16\times 16$.
@@ 5.2 Training Performance and Scaling
To evaluate the practical utility of our approach, we integrated the proposed matrix exponential method into a generative framework. Glow [37] is a coupling layer approach to learning invertible transformations. We replaced the original implementation (expm_flow) with our optimized function (expm_flow_opt15) to measure its impact on computational efficiency during the training phase. All models were trained for 50 epochs across the three aforementioned benchmark datasets CIFAR-10, ImageNet32, and ImageNet64. As detailed in Table 6, the proposed method demonstrates increased computational efficiency across all tested configurations. While the minimum observed speed-up was $3.91$ for ImageNet64, the most significant gain occurred with ImageNet32, where training was nearly 10 times faster ($9.74$) than the baseline. This dramatic reduction in training time, specifically the shift from 32304 seconds to 3317 seconds per epoch on ImageNet32, shows that the computational bottleneck previously associated with matrix exponentials in flow-based models can be effectively mitigated.
@@
expm_flow_opt15 time 443.4 3317 10875
@@
Table 6: Average training time per epoch (in seconds) for the matrix exponential version of Glow. Results compare the efficiency of our proposed expm_flow_opt15 function against the original expm_flow implementation across three standard datasets.
@@ 5.3 Inference and Sampling Performance
While the previous sections evaluated performance during the training phase, efficiency during inference is equally critical for generative models. Table 7 presents a comparison of the sampling times between expm_flow and expm_flow_opt15. In this context, sampling time refers to the duration required for the model to generate an image from the latent space. We evaluated two scenarios: generating a single image ($n=1$) and generating a batch of 128 images ($n=128$). The results demonstrate that the speedup provided by our method becomes significantly more pronounced as the batch size increases. While the performance is nearly identical for a single sample, the expm_flow_opt15 variant achieves a speedup factor of approximately $1.95$ for larger batches, effectively halving the inference latency.
@@
expm_flow_opt15 time 0.3836 0.4247
@@
Table 7: Average inference times (in seconds) for sampling a matrix exponential variant of Glow. We compare the original expm_flow against expm_flow_opt15 for single and batched samples, including the calculated speedup factor.
@@ 6 Conclusions
These results demonstrate that modern polynomial evaluation strategies allow for higher-order approximations at a lower computational cost, measured in terms of matrix multiplications ($M$), than traditional Paterson–Stockmeyer-based methods. Consequently, the proposed implementation offers a portable, library-independent solution for high-throughput applications.